A random sample of size 36 is taken from a population with mean µ = 29 and standard deviation σ = 5. What are the expected value and the standard deviation for the sampling distribution of the sample mean?

(a) 12.425 and 0.83, respectively

(b) 12.425 and 2.66, respectively

(c) 29 and 0.83, respectively

(d) 29 and 5, respectively
Answer #1

since expected value  for the sampling distribution of the sample mean =population mean=29

and standard deviation for the sampling distribution of the sample mean =σ /√n =5/√36=0.83

option C is correct

(c) 29 and 0.83, respectively